OpenStudy (anonymous):

Create a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -6) x+y=2 x-y=4

OpenStudy (anonymous):

easier to add them up and get \[2x=6\] then sole for \(x\)

OpenStudy (anonymous):

*solve

OpenStudy (anonymous):

x is 3 ?

OpenStudy (anonymous):

a graph is a long and time consuming way to solve this, but it will work yes \(x=3\) http://www.wolframalpha.com/input/?i=x%2By%3D2%2Cx-y%3D4

OpenStudy (anonymous):

and since \(x=3\) and \(x+y=2\) you have \(3+y=2\) and you can solve for \(y\)

OpenStudy (anonymous):

is y 5

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im so bad in alg 2 :(

OpenStudy (anonymous):

\[3+y=2\] subtract \(3\) from both sides to get \(y\)

OpenStudy (anonymous):

i think you added 3 but you need to subtract 3 a little practice is all you need, it is not like you are born knowing this

OpenStudy (anonymous):

Oh yeah i know how to do that, i was good ion alg 1 & geometry when i was in public school

OpenStudy (anonymous):

so y is -1

OpenStudy (anonymous):

yes, \(y=-1\)

OpenStudy (anonymous):

you can check that this is right, because \[3-1=2\\3+1=4\]

OpenStudy (anonymous):

ok i got you, & so i write the answer like this? (3,-1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome! Thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I got it wrong!?

OpenStudy (anonymous):

no, it is right, the solution is \((3,-1)\)

OpenStudy (anonymous):

It told me it was wrong..

OpenStudy (anonymous):

wait i typed it on the wrong one :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

damn on line systems all ways messing with you

OpenStudy (anonymous):

I know! If i wouldnt have done that i woulda passed that damn quiz!

OpenStudy (anonymous):

what? you have to do it again??

OpenStudy (anonymous):

Yea i gotta retake it, im doing online school :/

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Use a graph and a table to solve each system. Check your answer. y+x=5 3x-5y=-1

OpenStudy (anonymous):

can u explain step by step how to do these?

OpenStudy (anonymous):

do you have to use a graph? or can we just solve it?

OpenStudy (anonymous):

it would be much easier and quicker to solve using algebra rather than a graph

OpenStudy (anonymous):

yea we can just sole its multi choice no graph stuff

OpenStudy (anonymous):

ok then lets knock it out

OpenStudy (anonymous):

\[y+x=5\\ 3x-5y=-1\] first equation is the same as \(y=5-x\) so we can replace the \(y\) in the second equation by \(5-x\) with careful use a parentheses and get \[3x-5(5-x)=-1\]

OpenStudy (anonymous):

this has only one variable, so we can solve for \(x\) via \[3x-5(5-x)=-1\\ 3x-25+5x=-1\\ 8x-25=-1\]

OpenStudy (anonymous):

now add \(25\) to both sides and get \[8x=24\] and solve for \(x\) by dividing by \(8\) to get \(x=3\)

OpenStudy (anonymous):

now since \(x=3\) and \(x+y=5\) we get \(3+y=5\) making \(y=2\) and the solution is \((3,2)\)

OpenStudy (anonymous):

more or less clear?

OpenStudy (anonymous):

still so confusing but i can use that as a guide! Thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Do you know this? I have got it wrong 2 times already! Classify the system and determine the number of solutions. Check all that apply. 7x + y = 13 28x + 4y = -12 infinitely many solutions inconsistent consistent no solution

OpenStudy (anonymous):

this is NO Solution for this also known as "inconsistent"

OpenStudy (anonymous):

i think you are supposed to select them both

OpenStudy (anonymous):

\[7x + y = 13 \\28x + 4y = -12\]if yo multiply the first equation by \(4\) you get \[28x + 4y = 52\\ 28x + 4y = -12\] and there is no way to make \(52=-12\)

OpenStudy (anonymous):

what about this one?? Classify the system and determine the number of solutions. (Check all that apply.) $7x - y = -11$ $3y = 21x + 33$ dependent inconsistent infinitely many solutions no solution consistent independent

OpenStudy (anonymous):

what's with the dollar signs?

OpenStudy (anonymous):

Oh it just put that in there ignore those, i just copied it from my page

OpenStudy (anonymous):

ok well they are the same line, so one answer is "infinitely many solutions" also i believe this can be called "consistent"

OpenStudy (anonymous):

oh, also add "dependent"

OpenStudy (anonymous):

Thank you so so much!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

did you pass the quiz?

OpenStudy (anonymous):

im trying to do these last 2 by going with what u explained im so freaking lost!

OpenStudy (anonymous):

i will help you if you like

OpenStudy (anonymous):

Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -6) x - y = -1 4x - 2y = 2

OpenStudy (anonymous):

See i had a baby and thats the only reason im finishing my senior year online. & its so hard with the math, i got behind so i hvae 3 units to do by the 9th!

OpenStudy (anonymous):

\[x - y = -1 \\4x - 2y = 2\] multiply the first equation by \(-2\) and get \[-2x+2y=2\\ 4x-2y=2\]

OpenStudy (anonymous):

now add them up and get \(2x=4\) and so \(x=2\)

OpenStudy (anonymous):

since \(x=2\) and \(x-y=-1\) you have \(2-y=-1\) making \(y=3\) and the solution is \((2,3)\) how old is your baby?

OpenStudy (anonymous):

That was much easier to understand, now i just gotta get it to stick with me, & he will be 7 months on the 6th

OpenStudy (anonymous):

must keep you pretty busy i am sure hard to do math, on line or otherwise

OpenStudy (anonymous):

you done?

OpenStudy (anonymous):

Yes i passed that time!!

OpenStudy (anonymous):

whew!!

OpenStudy (anonymous):

good luck with the rest of the class

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw